# FlashKAT: Understanding and Addressing Performance Bottlenecks in the Kolmogorov-Arnold Transformer

### v2

@@ 6 Conclusion
The successes of the KAN, due to its increased expressivity and interpretability, have made it a popular alternative to the MLP. However, it has yet to achieve widespread application in large-scale NLP and computer vision tasks due to its increased parameters, computational cost, lack of GPU hardware optimization, and training instability. The KAT has resolved many of the KAN’s shortcomings with the introduction of GR-KAN, but it still suffers from slow training speeds. In this work, we conduct experiments to understand the source of the slowdown, and we identifiy that the slow training speeds are a byproduct of memory stalls in the backward pass, a discovery that eluded previous works (yang2024kolmogorov; molina2019pad).
To address this bottleneck, we propose a novel restructuring of the group-wise rational function that allows efficient gradient accumulation by minimizing atomic add usage. The resulting FlashKAT achieves an 86.5x speedup compared to KAT, while reducing gradient rounding errors. The application of our work both enables wider adoption and expedites research on the KAT.
Limitations. A main drawback preventing the use of KAN in the Transformer is the increased computation of the KAN. With GR-KAN, the KAT successfully reduced the FLOPs to nearly match the speed of an MLP on the forward pass, but it is still orders of magnitude slower on the backward pass, making training a computational burden. Our proposed FlashKAT, with a significant speedup, has mostly filled the performance gap, but is still currently about 20% slower than a standard Transformer. Even so, our work is an important step in closing the gap between the KAT and standard Transformer in terms of training speed, motivating future work to narrow the gap even further.

### v3

@@ 6 Conclusion
Although KAN has achieved success, it has yet to see widespread adoption in large-scale NLP and computer vision tasks due to its increased parameter count, higher computational cost, lack of GPU hardware optimization, and training instability. KAT has addressed many of KAN’s shortcomings with the introduction of GR-KAN, but it still suffers from slow training. In this work, we conduct experiments to understand the source of the slowdown and identify that slow training speeds are a byproduct of memory stalls in the backward pass, a discovery that eluded prior work (Yang and Wang 2024; Molina et al. 2019; Delfosse et al. 2020, 2021). To address this bottleneck, we propose a novel restructuring of the group-wise rational function that allows efficient gradient accumulation by minimizing atomic add usage. The resulting FlashKAT achieves an 86.5x speedup over KAT, while reducing gradient rounding errors. The application of our work both enables wider adoption and expedites research on KAT.
Limitations. A main drawback preventing the use of KAN in the Transformer is KAN’s increased computational cost. With GR-KAN, KAT successfully reduced FLOPs, allowing it to nearly match the speed of a standard Transformer on the forward pass. Still, it remains orders of magnitude slower on the backward pass, making training computationally burdensome. Our proposed FlashKAT, with a significant speedup, has largely closed the performance gap, but is still about 25% slower than a standard Transformer. Even so, our work is an important step toward closing the gap between the KAT and standard Transformer in training speed, motivating future work to narrow it further.